OpenStudy (morganwarren):

PLEASE help A 25 gram sample of a substance that's used for drug research has a k-value of 0.1229 Find the substance's half-life in days.

OpenStudy (anonymous):

set \[\huge e^{-0.1229t}=0.5\] and solve for \(t\) in two steps

OpenStudy (morganwarren):

i have literally no clue what i'm doing

OpenStudy (anonymous):

ok then lets go slow

OpenStudy (anonymous):

first off, the 25 grams has nothing to do with anything you are looking for a half life, the time it takes for the substance to be reduces to one half its original amount

OpenStudy (anonymous):

*reduced

OpenStudy (anonymous):

second the "k value" means the number that goes in the sky for your exponential decay function \[A(t)= A_0e^{-kt}\] you seen something like that before?

OpenStudy (morganwarren):

No not really I've just been guessing

OpenStudy (anonymous):

hmm then you can't really do this problem although it only takes two steps and a calculator

OpenStudy (wolf1728):

e^-.1229*t = .5 Taking logs of both sides -.1229*t * Log (e) = Log (.5) e= 2.71828 -.1229*t * 0.4342944818 = -0.3010299957 t = -0.3010299957 / (-.1229 * 0.4342944818) If you have a calculator you can calculate the answer.

OpenStudy (morganwarren):

5.6?

OpenStudy (wolf1728):

Yes, 5.6 is correct